OpenStudy (anonymous):

1.) Divers looking for a sunken ship have defined the search area as a triangle with adjacent sides of length 2.75 miles and 1.32 miles. The angle between the sides of the triangle is 35°. To the nearest hundredth, find the search area My answer: 1.04 mi squared. Is that correct?

OpenStudy (anonymous):

2.) Name the major arc and find its measure. http://assets.openstudy.com/updates/attachments/4fc429ece4b0964abc86c269-kaitlinlikesmusic-1338255864730-arc.png My answer: ABD310 degrees. Is that correct?

OpenStudy (rhr12):

second one is correct which is about the major arc

OpenStudy (rhr12):

first one is also correct

OpenStudy (rhr12):

you did a great job

OpenStudy (rhr12):

A = (1/2)*ab*sin(C) A = (1/2)*2.75*1.32*sin(35) A =~ 1.04 square miles

OpenStudy (rhr12):

this is about the first one

OpenStudy (rhr12):

i don't think i have to explain it